Mode p+q mode = mode p-q mode,then what is the angle between vector p and q

Grade:11

1 Answers

Arun
5 years ago

If we solve the question by vectors,

Magnitude of P+Q=magnitude of P-Q

P^2+Q^2+2PQcosx=P^2+Q^2–2PQcosx where x is the angle between the two quantities

Therefore,

4PQcosx=0

Cosx=0

Therefore x=90 degrees or pi/2 radian